# Statistical Demography Meets Ministry of Health: The Case of the Family Planning Estimation Tool

### v1

@@ 4.2 Process model: How do FP indicators vary over time and differ across populations?
Here, $\theta^{(l-1)}_{\gamma^{\prime}}$ is the parameter value at one level higher in the hierarchy for the region $\gamma^{\prime}$ that contains unit $\gamma$. At the top level of the hierarchy, $l=1$, only a single parameter is estimated, representing an intercept that applies to all lower level geographic units. The parameter $\sigma_{\theta}^{(l)}$ for $l>1$ describes variability of values of the parameter $\theta$ across geographic units at level $l$ of the hierarchy.
Illustrative findings
Illustrative results of model fits are shown in Figure 5 for Angola, Bolivia, and Burundi. We see that the model can produce estimates and forecasts for all years. Uncertainty associated with FP indicators increases with the forecast horizon.
Figure 5: Data, estimates, and forecasts from 1970 until 2030 for FP indicators for married women in Angola, Bolivia, and Burundi. Graphs show mCPR, unmet need for modern methods, demand, and demand satisfied with modern methods. Survey data are shown in colored dots with 95% confidence intervals that reflect total error variance. FPET estimates are shown in black, with grey shaded areas representing 90% credible intervals.
4.3 Data model for survey data
@@
The model is fitted to survey data on FP indicators, while accounting for different sources of error. First, all observations are subject to sampling error arising from the survey design. In addition, other non-systematic errors may occur which we will refer to as measurement errors. Measurement errors may occur due to errors introduced during data collection or compilation. The extent of such errors may vary based on survey characteristics. For example for mCPR estimation, some survey programs have been found to provide more variable measurements of mCPR than others (Alkema et al. [2013]). Extreme errors may also occur, as we saw in the mCPR example for Burundi in Figure 1, resulting in outlying observations due to measurement error. Finally, errors may occur if the sampled population differs from the population of interest. For mCPR estimation, such errors are typically absent from recently collected data by international programs but may be present in older data or data from national or other survey programs. Examples include variation in age groups (Alkema et al. [2013], Kantorová et al. [2020]). When using survey data, we need to account for data errors, to produce estimates that are aligned with observations that are very certain and deemed to be close to the true value, while smoothing over outlying data points that are very uncertain.
@@
The survey data model specifies the relation between the observed survey data and the FP indicators that are estimated. In FPET, the normal-with-optional-shrinkage (NOS) class of data models is used to account for the types of errors that may occur (Alkema et al. [2024a]). This class of data models accounts for sampling errors and differences in observational uncertainty based on survey characteristics. In addition, the NOS data model uses shrinkage priors to produce estimates that are robust to outlying observations. The data model class also accounts for settings where errors may be correlated across observations, as is the case for FP with observations from PMA data series.
@@
Details on the the NOS data model and its use in FPET are provided in Alkema et al. [2024a]; we summarize the main set up here. We consider the NOS data model for the logit-transform of two observed proportions: mCPR and unmet need for modern methods out of non-modern-users. With slight misuse of notation for ease of exposition, let $y_{i}$ refer to one observed proportion and let $\phi_{c[i],t[i]}$ refer to the true latent indicator for the corresponding country-year. For example, let $y_{i}$ refer to observed mCPR and $\phi_{c[i],t[i]}$ to true mCPR. The total error associated with $y_{i}$, on the logit-transformed scale, is decomposed into type-specific errors:
@@
For the sampling error $E_{i}^{(\mathsf{samp})}$, the associated variance is $\sigma_{i}^{(\mathsf{samp})2}=s_{i}^{2}$, where $s_{i}$ is the sampling error on the logit-scale. Source-type errors capture that errors from some programs may be larger than errors from other programs: $\sigma_{i}^{(\mathsf{source})}=\tilde{\sigma}_{d[i]}^{(\mathsf{source})}$, where $d[i]\in\{1,\dots,D\}$ indexes the survey source of observation $i$, i.e., DHS, MICS, PMA, National survey, or other survey. For mCPR estimation, we assign source errors to observations from all sources except those from DHS, i.e., $\mathcal{D}$ excludes DHS. The term $E_{i}^{(\mathsf{char})}$ with variance $\sigma^{(\mathsf{char})2}$ is added if a survey did not capture the entire population of interest, for example, if ever-married women were interviewed as compared to married women. The subset of observation indices for which the population surveyed differs from that of interest, $\mathcal{C}$, is obtained from survey meta data.
The outlier error variance arises from a shrinkage prior on the error term (Piironen and Vehtari [2017]) that shrinks outlier errors towards zero for most observations, only allowing large errors when the observation is extremely different from the trend. In this set up, the standard deviation term $\sigma_{i}^{(\mathsf{outlier})}$ is parameterized using a scale parameter $\tau$ that is shared across observations, and an observation-specific scale parameter $\gamma_{i}$ that allows individual errors to escape from shrinkage, $\sigma_{i}^{(\mathsf{outlier})}=\tau\tilde{\gamma}_{i}=\sqrt{\frac{\tau^{2}\vartheta^{2}\gamma_{i}^{2}}{\vartheta^{2}+\tau^{2}\gamma_{i}^{2}}}$. The local scale parameter $\gamma_{i}$, which is assigned a heavy-tailed half-Cauchy prior, $\gamma_{i}\sim C^{+}(0,1)$. Observation indices $\mathcal{I}$ refer to the set of data points that may be possibly outlying, determined by a data classification algorithm [Alkema et al., 2024a].
@@
with $\sigma_{i}^{2}=s_{i}^{2}+\mathbb{I}(d[i]\in\mathcal{D})\sigma_{i}^{(\mathsf{source})2}+\mathbb{I}(i\in\mathcal{C})\sigma_{i}^{(\mathsf{char})2}+\mathbb{I}(i\in\mathcal{I})\sigma_{i}^{(\mathsf{outlier})2}$. We assume observations are conditionally independent except for data collected by the PMA program, as results from consecutive PMA surveys may be correlated because of repeated sampling of individuals across survey rounds. For PMA data series for a country-marital group, we use the multivariate normal specification:
@@
Illustrative findings
Survey data and FPET fits are shows for Burundi, Ethiopia, and Niger in Figure 6. The fit for Burundi shows that based on the data outlier error set up, the estimates smooth over the second-last data point. In Ethiopia, FPET estimates are close to the most recent DHS, accounting for correlation in the PMA data series, that trends lower. FPET estimates suggest some uncertainty in recent years in Niger given conflicting data from PMA and a national survey. In this instance, if there is information on data quality, users may label a more recent survey as non-outlying to produce estimates based on that type of input (see discussion).
Figure 6: Data and estimates for mCPR among married women for Burundi, Ethiopia, and Niger. Data are shown in colored dots with 95% confidence intervals that reflect total error variance. FPET estimates are shown in black, with grey shaded areas representing 90% credible intervals.
4.4 Data model for service statistics data
@@
Service statistics data are used in FPET to inform mCPR estimates. There are four types of FP service statistics data: the number of contraceptive commodities distributed to clients, the number of contraceptive commodities distributed to facilities, the number of times clients interacted with a provider for contraceptive services, and the number of current contraceptive users of any method. Track20 developed a tool to calculate an Estimated Modern Use (EMU) indicator from these different types of service statistics data [Track20, 2020, 2024]. EMU estimates have been used in FPET to inform mCPR estimates in the absence of recent survey data [Cahill et al., 2018]. Here we summarize the approach used in the most recent version of FPET. Additional details can be found in Mooney et al. [2024a].
@@
$\Delta EMU_{d,c,t}|\sigma_{EMU,d,c,t}\sim N(\Delta\mathrm{logit}(\phi_{c,t}),\sigma_{EMU,d,c,t}^{2}),$ (9)
where $\Delta EMU_{d,c,t}$ refers to the observed change in EMU for data type $d$, country $c$, and year $t$. Error variance $\sigma_{EMU,d,c,t}^{2}$ is given by
$\sigma_{EMU,d,c,t}^{2}=s_{EMU,d,c,t}^{2}+\sigma_{EMU,d,c}^{2}$ (10)
with $s_{EMU,d,c,t}^{2}$ the pre-calculated variance based on uncertainty in inputs and adjustments, and $\sigma_{EMU,d,c}^{2}$ a country-type specific variance. We obtain $s_{EMU,d,c,t}^{2}$ based on a Monte Carlo simulation, using uncertainty in inputs associated with the adjustments of EMUs to capture the private share (Mooney et al. [2024b]). The country-type variance is estimated hierarchically:
@@
Illustrative findings
The effect of including EMUs is illustrated in a case study in Figure 7. The top row shows information on mCPR (left) and rates of change in mCPR (right), based on EMUs (red) and survey-based estimates of mCPR (black). The vertical lines indicate the 95% confidence intervals for mCPR based on individual EMUs and their respective $s_{EMU,d,c,t}$ alone. The EMU-based rates of change suggest a stall after the most recent survey year, with rates of change that are lower than survey-based predictions. Plots in the bottom row illustrate the effect of including the EMUs in FPET estimates: the stall in EMUs past the most recent survey results in a slower estimated mCPR growth, as compared to survey-only estimates.
Figure 7: Illustration of the effect of including EMUs when estimating mCPR. The top row shows observed EMUs (red) for the level (left) and rate of change (right). The vertical lines indicate the 95% confidence intervals based on $s_{EMU,d,c,t}$ alone. Survey-based estimates of mCPR are added in black. The blue dashed line indicates the most recent survey year. The top row illustrates that the EMU-based rates of change suggest a stall after the most recent survey year, with rates of change that are lower than survey-based predictions. Plots in bottom row illustrate the effect of including the EMUs in FPET estimates: the stall in EMUs past the most recent survey results in a slower mCPR growth, as compared to survey-only estimates.
4.5 Computation and sequential model fitting
The model was implemented in the Stan programming language [Stan Development Team, 2024, Gabry et al., 2024], and all analyses were conducted using the R statistical computing environment [R Core Team, 2022]. Code is available in an open-source R package [Alkema et al., 2024b].
In a flexible process model that combines short- and long-term trends, it may be challenging to identify short-term acceleration or deceleration versus a longer term change in rate of change. This challenge holds true for the transition model used in FPET: rates of change combine longer term trends, captured by the transition function, with shorter-term accelerations or decelerations that are captured by the smoothing term. To overcome possible issues, we carried out sequential fitting to first estimate long-terms trends, followed by the assessment of short-term fluctuations. To estimate long-term trends, we fit a model with transition functions without smoothing components to the global data base. This step results in the estimation of subcluster transition functions that represent long-term trends. We use the estimated subcluster transition functions from this step as inputs in subsequent model fitting steps, i.e., to estimate parameters that govern short-term trends, as well as data model parameters.
The main utility of FPET is as a tool for countries to produce estimates for a particular population of interest. We provide this implementation through fixing non-country-specific parameters at point estimates from the global fit. The resulting “local model” can be fit to data for just one country. A web-based application of the local model with a user-friendly interface is available at https://fpet.track20.org/. This tool allows for user inputs in instances where substantive knowledge conflicts with automated categorization of observations into being possibly outlying or not. Specifically, if an observation is outlying but is deemed to capture the truth, the user has the option to change the possible outlier status and hence remove the outlier error term from the data model. Vice versa, when an observation from the reference category is deemed subject to data quality concerns, the error term for possible outliers can be added.

### v2

@@ 3.2.2 Process model: How do FP indicators vary over time and differ across populations?
Here, $\theta^{(l-1)}_{\gamma^{\prime}}$ is the parameter value at one level higher in the hierarchy for the region $\gamma^{\prime}$ that contains unit $\gamma$. At the top level of the hierarchy, $l=0$, only a single parameter is estimated, representing an intercept that applies to all lower level geographic units. The parameter $\sigma_{l,\theta}$ for $l>1$ describes variability of values of the parameter $\theta$ across geographic units at level $l$ of the hierarchy. Priors are defined for the intercept and the variance terms:
$\theta^{(0)}$ $\sim N\left(s_{0,\theta}\cdot m_{0,\theta},s_{0,\theta}^{2}\right),$
$\sigma_{l,\theta}$ $\sim N^{+}\left(0,w_{0,\theta}^{2}\right),$
where $s_{0,\theta}\cdot m_{0,\theta}$ and $s_{0,\theta}$ refer to the prior mean and standard deviation of $\theta^{(\mathrm{global})}$, and $w_{0,\theta}$ refers to the prior standard deviation for $\sigma_{l,\theta}$.
Specific settings for all parameters are given in the Appendix. As an example, we consider the parameter for the level of transformed demand among married women in reference year $t^{*}$, given by $\Omega_{c}$. No transformation is used for this parameter because $\Omega$ is already unconstrained. The hierarchical levels are given by world, cluster, subcluster, and country. The model set up is as follows, naming the hierarchical levels:
$\Omega_{c}^{(\mathrm{country})}\mid\Omega_{s[c]}^{(\mathrm{subcluster})},\sigma_{\mathrm{country},\Omega}^{2}$ $\sim N(\Omega_{s[c]}^{(\mathrm{subcluster})},\sigma_{\mathrm{country},\Omega}^{2}),$
$\Omega_{s}^{(\mathrm{subcluster})}\mid\Omega_{w[s]}^{(\mathrm{cluster})},\sigma_{\mathrm{subcluster},\Omega}^{2}$ $\sim N(\Omega_{w[s]}^{(\mathrm{cluster})},\sigma_{\mathrm{subcluster},\Omega}^{2}),$
$\Omega_{w}^{(\mathrm{cluster})}\mid\Omega^{(\mathrm{global})},\sigma_{\mathrm{cluster},\Omega}^{2}$ $\sim N(\Omega^{(\mathrm{global})},\sigma_{\mathrm{cluster},\Omega}^{2}),$
$\sigma_{l,\Omega}$ $\sim N^{+}(0,2^{2}),\text{ for }l=1,2,3$
$\Omega^{(\mathrm{global})}$ $\sim N(0,3^{2}).$
3.2.3 Data model for survey data
The survey data model specifies the relation between the observed survey data and the FP indicators that are estimated, such that the overall model can be fitted to the survey data while accounting for different sources of error.
The survey data base contains observations on the compositional vector of FP proportions, i.e., contraceptive use of modern and traditional methods, and unmet need for any method (see Section 2.2). For model fitting, we consider three proportions: mCPR, the ratio of unmet need for modern methods out of non-modern-users, and traditional users out of those with an unmet need for modern methods.
@@ 3.2.3 Data model for survey data
All observed proportions may be subject to various errors. First, all observations are subject to sampling error arising from the survey design. In addition, other non-systematic errors may occur which we will refer to as measurement errors. Measurement errors may occur due to errors introduced during data collection or compilation. The extent of such errors may vary based on survey characteristics. For example for mCPR estimation, some survey programs have been found to provide more variable measurements of mCPR than others (alkema2013mcpr). Extreme errors may also occur, as we saw in the mCPR example for Burundi in Figure 1, resulting in outlying observations due to measurement error. Finally, errors may occur if the sampled population differs from the population of interest. For mCPR estimation, such errors are typically absent from recently collected data by international programs but may be present in older data or data from national or other survey programs. Examples include variation in age groups (alkema2013mcpr, kantorova_estimating_2020). When using survey data, we need to account for data errors, to produce estimates that are aligned with observations that are very certain and deemed to be close to the true value, while smoothing over outlying data points that are very uncertain.
@@
To model the relation between FP indicators and an observed proportion, accounting for possible errors in the observed proportion, we use the normal-with-optional-shrinkage (NOS) class of data models (alkema_2024dm). This class of data models accounts for different types of errors that may occur. These errors include sampling errors and differences in observational uncertainty based on survey characteristics. In addition, the NOS data model uses shrinkage priors to produce estimates that are robust to outlying observations.
We summarize the main set up here for one proportion, i.e., mCPR. Additional details are given in the Appendix (Section 7.2.2).
@@
With slight misuse of notation for ease of exposition, let $y_{i}$ refer to one observed proportion and let $\phi_{c[i],t[i]}$ refer to the true latent indicator for the corresponding country-year. For example, let $y_{i}$ refer to observed mCPR and $\phi_{c[i],t[i]}$ to true mCPR. The total error associated with $y_{i}$, on the logit-transformed scale, is decomposed into type-specific errors:
@@
For the sampling error $E_{i}^{(\mathsf{samp})}$, the associated variance is $\sigma_{i}^{(\mathsf{samp})2}=s_{i}^{2}$, where $s_{i}$ is the sampling error on the logit-scale. Source-type errors capture that errors from some programs may be larger than errors from other programs: $\sigma_{i}^{(\mathsf{source})}=\tilde{\sigma}_{d[i]}^{(\mathsf{source})}$, where $d[i]\in\{1,\dots,D\}$ indexes the survey source of observation $i$, i.e., DHS, MICS, PMA, National survey, or other survey. The term $E_{i}^{(\mathsf{char})}$ with variance $\sigma^{(\mathsf{char})2}$ is added if a survey did not capture the entire population of interest, for example, if ever-married women were interviewed as compared to married women. The subset of observation indices for which the population surveyed differs from that of interest, $\mathcal{C}$, is obtained from survey meta data.
The final error term, referred to as outlier error, is designed to handle detection of outlier observations and smoothing over them. Outlier detection and smoothing is accomplished by assigning shrinkage priors to the outlier error terms (piironen2017horseshoe), to shrink outlier errors towards zero for most observations, only allowing large errors when the observation is extremely different from the trend. The outlier error standard deviation term $\sigma_{i}^{(\mathsf{outlier})}$ is parameterized using a scale parameter $\tau$ that is shared across observations, and an observation-specific scale parameter $\gamma_{i}$, $\sigma_{i}^{(\mathsf{outlier})}=\sqrt{\frac{\tau^{2}\vartheta^{2}\gamma_{i}^{2}}{\vartheta^{2}+\tau^{2}\gamma_{i}^{2}}}$. The local scale parameter $\gamma_{i}$ is assigned a heavy-tailed half-Cauchy prior, $\gamma_{i}\sim C^{+}(0,1)$, to allow individual errors to escape from shrinkage to zero.
@@
with $\sigma_{i}^{2}=s_{i}^{2}+\mathbb{I}(d[i]\in\mathcal{D})\sigma_{i}^{(\mathsf{source})2}+\mathbb{I}(i\in\mathcal{C})\sigma^{(\mathsf{char})2}+\mathbb{I}(i\in\mathcal{I})\sigma_{i}^{(\mathsf{outlier})2}$. We assume observations are conditionally independent except for data collected by the PMA program, as results from consecutive PMA surveys may be correlated because of repeated sampling of individuals across survey rounds. For PMA data series for a country-marital group, we use the multivariate normal specification:
@@
When to (not) include outlier error terms?
Outlier error terms are included only for observations with indices $\mathcal{I}$. This subset is introduced to avoid automatically smoothing over high-quality observations that are outlying but are deemed to capture true fluctuations in FP indicators.
A data classification algorithm is used to produce a set of reference survey data points that are, by default, not deemed to be outlying, as follows. For each country, we first label observations as possibly outlying if there are data quality concerns or other documented substantive information to suggest they may be subject to substantial error. For example, DHS observations collected before 1990 may be outlying because of possible misclassification of women who were pregnant or post-partum. A survey may be classified as being possibly outlying if issues with the survey process have been documented. Subsequently, for each country, we check if we can identify a reference source category among the left over points, which, as the name suggests, provides a reference in the estimation exercise and is assumed to be not subject to outlier errors. This category is given by DHS data, if available. If no DHS data are available, data from national surveys or other survey programs are used, with the choice based on the source type with the most observations after 1990 in that country. If the algorithm does not result in a reference category, i.e., if there are no non-possibly-outlying observations that are from DHS, national, or other surveys, then the population does not have a reference category and all observations are possibly outlying.
The default classification of possibly outlying points is used in global model fitting for estimation of long term trends. In local model fitting, users may update the classification to reflect external information about data quality.
3.2.4 Global model fitting and computation
Global model fitting refers to estimating all model parameters of the combined process and survey data model, as summarized in Figure 4. Global model fitting occurs separately by marital group, as summarized in Figure 3.
To estimate mCPR and unmet need for modern methods, per marital group, we use two steps. In step 1A, we estimate global and world region parameters governing long-term trends in demand and demand satisfied. In step 1B, parameters governing short-term fluctuations and data quality are estimated. The motivation for fitting sequentially in two steps is to avoid possible issues identifying long-term trends. Details are provided in Appendix Section 7.2.3.
The global model was implemented in the Stan programming language [stan2023, cmdstanr2022], and all analyses were conducted using the R statistical computing environment [r2022]. Code is available in the open-source R package fpet2, see fpetrpackage. The package’s website contains detailed instructions for global model fitting, see https://alkemalab.github.io/fpet2/articles/workflow_globalnationalfitting.html.
3.3 Local modeling of FP indicators
The main utility of FPET is as a tool for countries to produce estimates for a particular population of interest, by fitting a model to data from that population alone. As summarized in Figure 3, we use what we refer to as “local models“: models that are derived from global (hierarchical) models but adapted such that they can be fitted to the data from one population alone.
For FPET, we provide an implementation of local models through fixing non-country-specific parameters at point estimates from the global fit. The set up follows the set up introduced in the localhierarchy R package [alkema_2025localhierarchy]. This set up is introduced in the next section. In local models, countries may want to use service statistics data to inform FP estimates. This is introduced in Section 3.3.2.
A web-based application of the local model with a user-friendly interface is available at https://fpet.track20.org/.
3.3.1 Local model specification and implementation
For local modeling, we use the specification and implementation provided in the localhierarchy R package [alkema_2025localhierarchy]. For national estimation using the FPET model, we start by fitting the global model as explained in the previous section, using a global data base and estimating all model parameters, including data quality parameters and means and variance of hierarchical distributions. In the derived local model (see Figure 7), parameters that are not country-specific are fixed using parameter estimates (posterior means) from the global model fit.
Local modeling is implemented in the R package fpet2 ([fpetrpackage]). The package uses functionality from localhierarchy to pass estimates of non-country-specific parameters from global models to local models. Local models are available for estimation of one marital group only, and for estimating both marital groups and calculating all-women estimates.
[…]
Figure 7: Overview of local model specification. The local model specification follows that of the global model (Figure 4), with parameters that are not country-specific fixed at the estimates obtained from the global model fit.
3.3.2 Use of service statistics data
@@
Service statistics data are used in FPET to inform mCPR estimates. There are four types of FP service statistics data: the number of contraceptive commodities distributed to clients, the number of contraceptive commodities distributed to facilities, the number of times clients interacted with a provider for contraceptive services, and the number of current contraceptive users of any method. Track20 developed a tool to calculate an Estimated Modern Use (EMU) indicator from these different types of service statistics data [bietsch_estimated_2025, emutrack20, ss2emu_tool]. EMU estimates have been used in FPET to inform mCPR estimates in the absence of recent survey data [cahill2018mcpr]. Here we summarize the approach used in the most recent version of FPET. Additional details can be found in mooney_emu.
@@
$\Delta EMU_{d,c,t}|\sigma_{EMU,d,c,t}\sim N(\Delta\phi_{c,t,m},\sigma_{EMU,d,c,t}^{2}),$ (8)
where $\Delta EMU_{d,c,t}$ refers to the observed change in EMU for data type $d$, country $c$, and year $t$. $\Delta\phi_{c,t,m}$ refers to the change in modern contraceptive use for the same country-period for marital group $m$, where $m$ refers to all women or married women. Error variance $\sigma_{EMU,d,c,t}^{2}$ is given by
$\sigma_{EMU,d,c,t}^{2}=s_{EMU,d,c,t}^{2}+\sigma_{EMU,d,c}^{2},$ (9)
with $s_{EMU,d,c,t}^{2}$ the pre-calculated variance based on uncertainty in inputs and adjustments, and $\sigma_{EMU,d,c}^{2}$ a country-type specific variance. We obtain $s_{EMU,d,c,t}^{2}$ based on a Monte Carlo simulation, using uncertainty in inputs associated with the adjustments of EMUs to capture the private share (mooney_emu1). The country-type variance is estimated hierarchically:
@@
4 Results
In this section, we first present illustrative findings that show country-specific fits to highlight modeling aspects. We then discuss results from a model validation exercise, followed by illustrative findings in local modeling when including service statistics data. Results for all countries are given in Appendix Figure 11.
4.1 Illustrative findings
To illustrate different contraceptive transitions and the effect of hierarchical modeling, model fits for married women in Bolivia, Burundi, Ethiopia, and Niger are shown in Figures 8 and 9. The figures illustrate how the model set up, using transition functions, allows for estimation and forecasting in different settings.
Results for Burundi and Bolivia illustrate how the model can produce estimates and forecasts for all years, even if data are limited for recent years. The uncertainty associated with FP indicators increases with the forecast horizon.
In Burundi, the most recent national survey indicates that mCPR is high, as compared to DHS observations in surrounding years. The FPET fit for Burundi treats the national survey observation as implausibly high and smooths over it, i.e., the model does not shrink the outlier error for the national survey in Burundi based on information in surrounding years. If substantive information on mCPR in Burundi would suggest that the national survey accurately captured a temporary uptake, the outlier error could be removed in a local model fit, to produce estimates that reflect that reality.
In Ethiopia, FPET estimates are close to the most recent DHS, accounting for correlation in the PMA data series, that trend lower.
Niger illustrates a setting with limited recent FP data: the absence of recent DHS surveys — which typically have smaller sampling errors and are considered less susceptible to non‑sampling errors — means analysts must rely on PMA and national surveys that carry greater uncertainty. FPET’s role in this context is to integrate those uncertain sources and produce estimates with uncertainty that accurately reflect the available evidence. Consequently, FPET estimates for Niger are substantially more uncertain for recent years because of these data limitations.
Figure 8: Data, estimates, and forecasts from 1970 until 2030 for FP indicators for married women in Bolivia and Burundi. Graphs show mCPR, unmet need for modern methods, demand, and demand satisfied with modern methods. Survey data are shown in colored dots with 95% confidence intervals that reflect total error variance. FPET estimates are shown in black, with grey shaded areas representing 90% credible intervals.
Figure 9: Data, estimates, and forecasts from 1970 until 2030 for FP indicators for married women in Burundi and Ethiopia. Graphs show mCPR, unmet need for modern methods, demand, and demand satisfied with modern methods. Survey data are shown in colored dots with 95% confidence intervals that reflect total error variance. FPET estimates are shown in black, with grey shaded areas representing 90% credible intervals.